Sabor Inc. is a medical testing laboratory that performs several tests and analyses for hospitals in the area. Four of the tests that they perform require the use of a specialized machine that can supply 14,000 hours per year. Information on the four lab tests follows:

Test A Test B Test C Test D
Charging rate $65 $51 $48 $32
Variable cost $25 $18 $13 $8
Machine hours 3 2 1 0.5


What is the contribution margin per unit of machine time for Test D?

a.$32

b.$20

c.$48

d.$8

e.$24
Answer #1

contribution margin=Charging rate-Variable cost

=(32-8)=$24 per unit

contribution margin per unit of machine time= contribution margin/Machine hours

=24/0.5

=$48.